# Learning Quantum Processes with Quantum Statistical Queries

### v1

@@ 3 Quantum Statistical Queries to Quantum Processes
Let :$\mathcal{E}\mathrel{\mathop{\ordinarycolon}}\mathbb{C}^{d}\rightarrow\mathbb{C}^{d}$ be a quantum process acting on a $d$-dimensional Hilbert space. A QPSQ learning algorithm has access to a quantum statistical query oracle QPStat of the process $\mathcal{E}$, which receives as input an observable $O\in\mathbb{C}^{d}\times\mathbb{C}^{d}$ satisfying $\mathinner{\!\left\lVert O\right\rVert}\leq 1$, a tolerance parameter $\tau\geq 0$, and using $poly(1/\tau)$ copies of a quantum state $\rho\in\mathcal{S}_{d}$, outputs a number $\alpha$ satisfying
@@
We denote the query as $\alpha\leftarrow\mathsf{QPStat}_{\mathcal{E}}(\rho,O,\tau)$. The output $\alpha$ acts as an estimate of the expectation value of $O$ on the state $\rho$ after evolution under $\mathcal{E}$ within absolute error $\tau$.
@@
We now discuss what it means for a QPSQ learning algorithm to be efficient. Any such learning algorithm aiming to learn some property of the process $\mathcal{E}$ must be a quantum polynomial time (QPT) algorithm, making at most polynomially many queries to the oracle QPStat. As the learner itself must provide the input states or their classical description to QPStat, it must be possible to efficiently prepare the required copies of these states. This is an important point for the physical justification of this model. If the learner was able to query QPStat with arbitrary quantum states that require exponential quantum computation for preparation, the learning model would no longer be physically justified. In the following definition for an efficient QPSQ learning algorithm, we only discuss the efficiency of the algorithm, not its correctness, allowing various notions of correctness depending on the desired property of $\mathcal{E}$ to be learned up for consideration.
@@
A QPSQ learning algorithm is called an efficient QPSQ learner if and only if it makes at most $\mathrm{poly}(\log(d))$ queries with tolerance at least $1/\mathrm{poly}(\log(d)$ to ${}_{\mathcal{E}}$ oracle, and runs in $\mathrm{poly}(\log(d))$ time.
@@
Along with the definition of Qstat, Arunachalam et. al.[9] presented algorithms for learning various concept classes in their QSQ model. The generalization of Qstat by QPStat implies that these algorithms also hold in our QPSQ learning model. Given QPStat oracle access to the process $\mathcal{E}_{c}$ associated with the target concept, and sufficient copies of the state $|\psi_{D}\rangle$, the required output from Qstat queries can be obtained using QPStat queries, and the remainder of the learning algorithms proceed identically. Thus, any concept class efficiently learnable in the QSQ model can be learned efficiently given QPSQ access to the unitary encoding of the target function instead.
@@ 4 Learning Arbitrary Processes from QPSQs
In this section, we introduce an algorithm for learning to predict the output of arbitrary quantum processes over a restricted class of observables. Our algorithm is very much inspired by the hybrid algorithm by Huang et al. in [46], which uses ideas from classical shadow tomography [48], to efficiently learn to predict the outcome of arbitrary quantum processes. In fact, our main observation was that the QPSQ model provides a more natural and general way of extracting data from a quantum process or a quantum experiment than the data extracted through the classical shadow queries, used in [46]. The classical machine learning part of our learning algorithm becomes very similar to the one in [46] with the difference being that instead of using the classical shadow approximation, we use the QPStat output to estimate the measurement outcome. This allows us to prove a rigorous performance guarantee for our learning algorithm with a modified query complexity. We discuss the applications of this learning algorithm later in section 6.
@@
for all $P\in\mathcal{P}_{n},|P|\leq k=\Theta(log(1/\epsilon))$ do
@@
     if  $(\frac{1}{3})^{P}>2\tilde{\epsilon}$ and $|x_{P}^{\prime\prime}(O)|>2.3^{|P|/2}\sqrt{\tilde{\epsilon}}\sum_{Q\mathrel{\mathop{\ordinarycolon}}a_{Q}\neq 0}|a_{Q}|$  then
@@
$h(\rho)\leftarrow\sum_{P\mathrel{\mathop{\ordinarycolon}}|P|\leq k}\hat{\alpha}_{P}(O)tr(P\rho)$
@@
$N=\Omega\left(\frac{log(n^{k+\kappa}/\delta)}{(\tilde{\epsilon}-\tau)^{2}}\right)$ (14)
@@
$\underset{\rho\sim\mathcal{D}}{\mathbb{E}}|h(\rho)-tr(O\mathcal{E}(\rho))|^{2}\leq\epsilon+\tilde{\epsilon}\max(\mathinner{\!\left\lVert O^{\prime}\right\rVert}^{2},1),$ (15)
@@
$x_{P}(O)=\underset{\rho\sim\mathcal{D}^{0}}{\mathbb{E}}tr(P\rho)tr(O\mathcal{E}(\rho))$ (16)
@@
$O^{\prime}=\sum_{P\in\{I,X,Y,Z\}^{\otimes n}\mathrel{\mathop{\ordinarycolon}}|P|\leq k}3^{|P|}x_{P}(O)P$ (17)
@@
$N=\Omega\left(\frac{log(n^{k+\kappa}/\delta)}{(\tilde{\epsilon}-\tau)^{2}}\right)$ (20)
@@
$\begin{split}Pr(|x^{\prime}_{P}(Q)-x_{P}(Q)|>\tilde{\epsilon}-\tau)&\leq 2exp\left(\frac{2N^{2}(\tilde{\epsilon}-\tau)^{2}}{\sum_{i=1}^{N}(2)^{2}}\right)\\ &=2\text{ }exp\left(\frac{-N(\tilde{\epsilon}-\tau)^{2}}{2}\right)\\ &\leq exp(-log(n^{k+\kappa}/\delta))\\ &=\frac{\delta}{n^{k+\kappa}}\end{split}$
@@
$\Pr[\forall P,Q\in\mathcal{P}_{n}\mathrel{\mathop{\ordinarycolon}}|x^{\prime}_{P}(Q)-x_{P}(Q)|<\tilde{\epsilon}-\tau]\geq 1-\delta$ (23)
@@
Analogous to our definition of $x_{P}^{\prime\prime}(Q)$, the proof in [46] begins by defining a variable $\hat{x}_{P}(Q)$. Their proof uses a bound similar to equation 25 with $\hat{x}_{P}(Q)$ instead. We refer to their work for the rest of the proof as it proceeds identically to theirs after this bound. ∎
@@
$\mathbb{E}[\alpha]=Tr(O\mathcal{E}(\rho_{in}))$ (26)
@@
Under the assumption on the output of QPStat shown in equation 26, the number of queries to QPStat of tolerance $\tau$ to achieve the error bound of the rigorous guarantee of Theorem 1 is given by
$N=\Omega\left(\frac{\tau^{2}log(n^{k+\kappa}/\delta)}{\tilde{\epsilon}^{2}}\right)$ (27)
@@
$tr(P\rho_{l})y_{l}(Q)\in[tr(P\rho_{l})tr(Q\mathcal{E}(\rho_{in}))-\tau,tr(P\rho_{l})tr(Q\mathcal{E}(\rho_{in}))+\tau],\forall l=1,2,...n$ (28)
@@
$\begin{split}Pr(|x_{P}^{\prime\prime}(Q)-x_{P}(Q)|>\tilde{\epsilon})&\leq 2\text{ }exp\left(\frac{-2N^{2}\tilde{\epsilon}^{2}}{\sum_{i=1}^{N}\left(2\tau)^{2}\right)}\right)\\ &=2\text{ }exp\left(\frac{-N\tilde{\epsilon}^{2}}{2\tau^{2}}\right)\\ &\leq exp(-log(n^{k+\kappa}/\delta))\\ &=\frac{\delta}{n^{k+\kappa}}\end{split}$ (29)
@@
$Pr(|x_{P}^{\prime\prime}(Q)-x_{P}(Q)|>\tilde{\epsilon})\leq\delta$ (30)

### v3

@@ 3 Quantum Statistical Queries to Quantum Processes
Let $\mathcal{E}:\mathbb{C}^{d}\rightarrow\mathbb{C}^{d}$ be a quantum process acting on a $d$-dimensional Hilbert space. A QPSQ learning algorithm has access to a quantum statistical query oracle QPStat of the process $\mathcal{E}$, which receives as input an observable $O\in\mathbb{C}^{d}\times\mathbb{C}^{d}$ satisfying $\norm{O}\leq 1$, a tolerance parameter $\tau\geq 0$, and using $\mathsf{poly}(1/\tau)$ copies of a quantum state $\rho\in\mathcal{S}_{d}$, outputs a number $\alpha$ satisfying
@@
We denote the query as $\alpha\leftarrow\mathsf{QPStat}_{\mathcal{E}}(\rho,O,\tau)$ [^1] When referring to unitaries, we will sometimes abuse the notation and use $\mathsf{QPStat}_{U}$ to refer to the oracle associated with the unitary channel $\rho\rightarrow U\rho U^{\dagger}$.. The output $\alpha$ acts as an estimate of the expectation value of $O$ on the state $\rho$ after evolution under $\mathcal{E}$ within absolute error $\tau$.
@@
We now discuss what it means for a QPSQ learning algorithm to be efficient. Any such learning algorithm aiming to learn some property of the process $\mathcal{E}$ must be a quantum polynomial time (QPT) algorithm, making at most polynomially many queries to the oracle QPStat. As the learner itself must provide the input states or their classical description to QPStat, it must be possible to efficiently prepare the required copies of these states. This is an important point for the physical justification of this model. If the learner was able to query QPStat with arbitrary quantum states that require exponential quantum computation for preparation, the learning model would no longer be physically justified. On a similar note, the observables that an efficient learner provides to the oracle must also be efficiently measurable. Then, in the following definition for an efficient QPSQ learning algorithm, we only discuss the efficiency of the algorithm, not its correctness, allowing various notions of correctness depending on the desired property of $\mathcal{E}$ to be learned up for consideration.
@@
A QPSQ learning algorithm is called an efficient QPSQ learner if it makes at most $\mathsf{poly}(\log(d))$ queries with tolerance at least $1/\mathsf{poly}(\log(d))$ to the QPStatE oracle with states preparable in $\mathsf{poly}(\log(d))$ time, and observables measurable up to precision $\tau$ in $\mathsf{poly}(\log(d),1/\tau)$ time, and runs in $\mathsf{poly}(\log(d))$ computational time.
@@
Along with the definition of QStat, Arunachalam et. al. [24] presented algorithms for learning various concept classes in their QSQ model. The generalization of QStat by QPStat implies that these algorithms also hold in our QPSQ learning model. Given QPStat oracle access to the process $\mathcal{E}_{c}$ associated with the target concept, and sufficient copies of the state $|\psi_{D}\rangle$, the required output from QStat queries can be obtained using QPStat queries, and the remainder of the learning algorithms proceed identically.
Remark 1.
Any concept class efficiently learnable in the QSQ model under a known distribution $D$ can be learned efficiently given QPSQ access to the unitary encoding of the target function instead if $|\psi_{D}\rangle$ can be prepared efficiently. Thus, by extending the algorithms from [24], one can show that there exist efficient QPSQ algorithms for learning parities, juntas and DNFs under the uniform distribution. The generalization also holds in the unknown distribution setting assuming access to copies of $|\psi_{D}\rangle$.
@@ 4 Learning arbitrary processes from QPSQs
In this section, we introduce an algorithm for learning to predict the output of arbitrary quantum processes over a restricted class of observables. Our algorithm is very much inspired by the hybrid algorithm by Huang et al. in [53], which uses ideas from classical shadow tomography [32], to efficiently learn to predict the outcome of arbitrary quantum processes. In some physical settings, one may only be able to access the statistics of a physical experiment rather than individual measurement outcomes. We show that even in such restricted settings, it is possible to achieve similar guarantees to the learning algorithm from [53] by adapting it to the QPSQ setting. The classical machine learning part of our learning algorithm is similar to the one in [53] with the difference being that instead of using the classical shadow estimates, we use the QPStat output to estimate the measurement outcome. This allows us to prove a rigorous performance guarantee for our learning algorithm with a modified query complexity. We discuss the applications of this learning algorithm later in section 7.
@@
for all $P\in\mathcal{P}_{n},|P|\leq k=\Theta(\log(1/\epsilon))$ do
@@
     if  $(\frac{1}{3})^{P}>2\tilde{\epsilon}$ and $|x_{P}^{\prime\prime}(O)|>2.3^{|P|/2}\sqrt{\tilde{\epsilon}}\sum_{Q:a_{Q}\neq 0}|a_{Q}|$  then
@@
$h(\rho)\leftarrow\sum_{P:|P|\leq k}\hat{\alpha}_{P}(O)Tr(P\rho)$
@@
$N=\Omega\left(\frac{\log(n^{k+\kappa}/\delta)}{(\tilde{\epsilon}-\tau)^{2}}\right)$ (18)
@@
$\underset{\rho\sim\mathcal{D}}{\mathbf{E}}|h(\rho)-Tr(O\mathcal{E}(\rho))|^{2}\leq\epsilon+\tilde{\epsilon}\max\left(\norm{O^{\prime}}^{2},1\right),$ (19)
@@
$x_{P}(O)=\underset{\rho\sim\mathcal{D}^{0}}{\mathbf{E}}Tr(P\rho)Tr(O\mathcal{E}(\rho))$ (20)
@@
$O^{\prime}=\sum_{P\in\{I,X,Y,Z\}^{\otimes n}:|P|\leq k}3^{|P|}x_{P}(O)P$ (21)
@@
$N=\Omega\left(\frac{\log(n^{k+\kappa}/\delta)}{(\tilde{\epsilon}-\tau)^{2}}\right)$ (24)
@@
$\begin{split}\mathbf{Pr}(|x^{\prime}_{P}(Q)-x_{P}(Q)|>\tilde{\epsilon}-\tau)&\leq 2\exp\left(\frac{2N^{2}(\tilde{\epsilon}-\tau)^{2}}{\sum_{i=1}^{N}(2)^{2}}\right)\\ &=2\exp\left(\frac{-N(\tilde{\epsilon}-\tau)^{2}}{2}\right)\\ &\leq\exp(-\log(n^{k+\kappa}/\delta))\\ &=\frac{\delta}{n^{k+\kappa}}\end{split}$
@@
$\mathbf{Pr}\left(\forall P,Q\in\mathcal{P}_{n}:|x^{\prime}_{P}(Q)-x_{P}(Q)|<\tilde{\epsilon}-\tau\right)\geq 1-\delta$ (27)
@@
Analogous to our definition of $x_{P}^{\prime\prime}(Q)$, the proof in [53] begins by defining a variable $\hat{x}_{P}(Q)$. Their proof uses a bound similar to equation 29 with $\hat{x}_{P}(Q)$ instead. We refer to their work for the rest of the proof as it proceeds identically to theirs from this point. ∎
@@
$\mathbf{E}[\alpha]=Tr(O\mathcal{E}(\rho_{in}))$ (30)
@@
Under the assumption of equation 30 on the output of QPStat, the number of queries to QPStat of tolerance $\tau$ to achieve the error bound of the rigorous guarantee of Theorem 1 is given by
$N=\Omega\left(\frac{\tau^{2}\log(n^{k+\kappa}/\delta)}{\tilde{\epsilon}^{2}}\right)$ (31)
@@
$Tr(P\rho_{l})y_{l}(Q)\in[Tr(P\rho_{l})Tr(Q\mathcal{E}(\rho_{in}))-\tau,Tr(P\rho_{l})Tr(Q\mathcal{E}(\rho_{in}))+\tau],\forall l\in[n]$ (32)
@@
$\begin{split}\mathbf{Pr}\left(|x_{P}^{\prime\prime}(Q)-x_{P}(Q)|>\tilde{\epsilon}\right)&\leq 2\exp\left(\frac{-2N^{2}\tilde{\epsilon}^{2}}{\sum_{i=1}^{N}\left(2\tau)^{2}\right)}\right)\\ &=2\exp\left(\frac{-N\tilde{\epsilon}^{2}}{2\tau^{2}}\right)\\ &\leq\exp(-\log(n^{k+\kappa}/\delta))\\ &=\frac{\delta}{n^{k+\kappa}}\end{split}$ (33)
@@
$\mathbf{Pr}\left(|x_{P}^{\prime\prime}(Q)-x_{P}(Q)|>\tilde{\epsilon}\right)\leq\delta$ (34)